Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (morsiani at Thu Jun 11 08:26:52 METDST 1998) Success

Westerbork (Tony Foley at Thu Jun 11 08:56:11 METDST 1998) Success
No known errors.

Noto (C.Trigilio at Thu Jun 11 09:53:11 METDST 1998) Success

Noto (C.Trigilio at Thu Jun 11 10:24:21 METDST 1998) Minor failures
Unfortunatelly the second sampler of the VLBA terminal had a failure and was not working. Recorded tracks are come from bbc 1,3,9,11; Lost signal from bbc 5,7,13.

Jodrell (Alastair Gunn at Thu Jun 11 11:35:20 METDST 1998) Success
No known problems.

Torun (K. Borkowski at Thu Jun 11 12:39:13 METDST 1998) Success
No known problems.

Onsala (A. Polatidis at Thu Jun 11 12:52:43 METDST 1998) Success

Calibration (Jean\-Francois Desmurs at Thu Jun 11 15:25:48 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may98 directory under the name ef005d.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Fri Jun 12 15:08:31 METDST 1998) Minor failures
Heavy rain around 0354

Effelsberg (dgraham at Tue Jun 16 16:18:37 METDST 1998) Failure
Unfortunately the operator did X-scan to point up, and found a big source 15 arc minutes away, so hea peaked up on that ! So we completely lost the program source. Sorry. This was only discovered today.